geerky42 (geerky42):

Find the range of the function: \(\Large y = \dfrac{x^2}{16x - 1} \)

geerky42 (geerky42):

@dpaInc

OpenStudy (anonymous):

since the degree of the numerator is 1 more than the degree of the denominator, you'll have a slant asymptote. this means the range will be a subset of all reals.

OpenStudy (anonymous):

are you familiar with finding local/relative extrema?

geerky42 (geerky42):

Kind of, I guess.

geerky42 (geerky42):

Ok, maybe not.

OpenStudy (anonymous):

ok find the derivative, y' = ???

OpenStudy (anonymous):

do you know about derivatives?

geerky42 (geerky42):

Applying quadrant rule: I got \(\Large y' = \dfrac{2x(16x - 1) - 16x^2 }{(16x - 1)^2} \).

OpenStudy (anonymous):

close... this is what i got: \(\large y'=\frac{2x(8x-1)}{(16x-1)^2} \)

geerky42 (geerky42):

Which can be reduced to \( \Large y' = \dfrac{16x^2 - 2x}{(16x^2-1)^2} \) It's the same. Now what?

OpenStudy (anonymous):

find the critical numbers....

OpenStudy (anonymous):

no it's not the same... look at what you gave as y' compared to what you have in the latter post... they're not the same.

OpenStudy (anonymous):

no matter, your last post for the derivative is correct now...

OpenStudy (anonymous):

now find critical numbers and determine whether they are local max or local mins...

OpenStudy (anonymous):

if you use the factored form of the derivative, you'll get x=0 and x=1/8. so the critical numbers will be x=0, x=1/8 by doing either first derivative test or second derivative test you'll find that x=0 is a local max; and x=1/8 is a local min.

OpenStudy (anonymous):

find the y values at these critical numbers and you'll get the boundaries for the range.